I would like to know how much force is applied on a stop sign held by a person when the wind is 80 km/hour against it perpendicularly? The stop sign is 16" x 16" (or 40.64 cm square).
1
Expert's answer
2012-04-03T11:37:38-0400
If the wind is blowing perpendicularly to the sign then the force can be& expressed as

F = S·V²,

where S is the area of the sign and V is the wind velocity. Also, we'll take into account that 40.64 cm² = 0.004064 m² and 80 km/hour = 80*1000/3600 m/s = 22.(2) m/s. So, the force is

F = 0.004064*1333.(3)² ≈ 2 N.